# Distributional Scaling Laws for Emergent Capabilities

### v2

@@ Abstract
In this paper, we explore the nature of sudden breakthroughs in language model performance at scale, which stands in contrast to smooth improvements governed by scaling laws. While advocates of “emergence” view abrupt performance gains as capabilities unlocking at specific scales, others have suggested that they are produced by thresholding effects and alleviated by continuous metrics. We propose that breakthroughs are instead driven by continuous changes in the probability distribution of training outcomes, particularly when performance is bimodally distributed across random seeds. In synthetic length generalization tasks, we show that different random seeds can produce either highly linear or emergent scaling trends. We reveal that sharp breakthroughs in metrics are produced by underlying continuous changes in their distribution across seeds. Furthermore, we provide a case study of inverse scaling and show that even as the probability of a successful run declines, the average performance of a successful run continues to increase monotonically. We validate our distributional scaling framework on realistic settings by measuring MMLU performance in LLM populations. These insights emphasize the role of random variation in the effect of scale on LLM capabilities.
Machine Learning, ICML
@@ 1 Introduction
On most benchmarks, language model (LM) performance is determined by a scaling law (Hestness et al., 2017; Rosenfeld et al., 2019; Kaplan et al., 2020) that varies smoothly with parameter size and overall training compute. There are, however, a number of celebrated exceptions in which performance abruptly improves on specific benchmarks (Srivastava et al., 2023).
@@
These sudden breakthroughs fuel one of the most heated debates in modern machine learning. On one side, advocates of emergence claim that performance abruptly improves at particular scales because those scales allow the LLM to acquire specific concepts that permit out-of-distribution generalization (Wei et al., 2022). On the other side, skeptics argue that these sudden improvements are a mirage driven by thresholding effects and alleviated by more appropriate continuous metrics—though a few breakthrough capabilities remain stubbornly emergent (Schaeffer et al., 2024). Here, we argue that such discontinuities are driven by continuous changes in the probability of a breakthrough at each scale.
@@
We posit that a breakthrough capability is often distinguished not by direct responses to scale, but by multimodal random variation—in other words, independent training runs form multiple clusters in their performance metrics. Such effects are currently undetected because scaling laws are usually determined by measuring a single training run at each scale, and resources are rarely committed to correct for random variation by reusing the same hyperparameter settings with different random seeds. Although random variation may be benign when model performance is measured in-distribution (Jordan, 2024), previous work suggests that out-of-distribution performance may vary widely across training runs (Zhou et al., 2024a, b), even at larger scales (Madaan et al., 2024). Since training numerous seeds becomes prohibitively expensive at large scales, we first study length generalization in small models trained on synthetic tasks. In similar settings, existing publications consider only a few training runs and report their summary statistics. By contrast, we characterize their full distribution, demonstrating that these breakthrough tasks exhibit multimodal variation. Our findings can be summarized as follows:
@@
Figure 1: Different random seeds produce different scaling trends. Scaling trends can be emergent or continuous for different seeds, even if all models train on the same data with the same hyperparameters. On the counting task, we show trends for random seeds with the highest breakthroughness (seed 93; top left) and linearity (seed 205; top right). We mark parameter counts immediately before and after seed 93’s emergence respectively as (a) and (b). Histograms illustrate the bimodal distribution of performance across all random seeds at scales (a) and (b), marking the positions of seeds 93 and 205. Breakthroughs occur when consecutive points represent different clusters; linear trends occur when each point is sampled from the same gradually shifting cluster.
@@
  • Breakthrough scaling results from bimodal performance distributions. Using length generalization as a case study, we demonstrate that different random seeds can exhibit either highly linear or highly emergent scale behavior. We connect these differences to the bimodal distribution of this compositional skill across random seeds, a property that materializes at many parameter scales. At these scales, “emergence” is a stochastic property. We observe these bimodal distributions around breakthrough scales both in small synthetic tasks and in LLM performance on MMLU.
@@
  • When a scale curve exhibits sudden, discontinuous improvement in a skill, the probability of a model learning that skill may be responding continuously to scale. By modeling the bimodal distribution as a mix of failure and success distributions, we illustrate that improvements with scale can come from changes in the probability of success or in the average performance of a successful model (Section 3). In response to the distributional nature of model performances, we propose to use Wassterstein distance to measure changes in performance distribution across model scale (Section 3.3). Using this metric, we show that the unlocking of model capacity at some small threshold in fact leads to a sudden emergence of bimodality in performance distribution from a unimodal one.
@@
  • Competing solutions can lead to either monotonic or U-shaped trends in emergence likelihood. We provide two synthetic tasks in which one has well-behaved scaling behavior and another exhibits U-shaped trends when increasing the width of the network (Section 3.4), a phenomenon also observed in LLM scaling (McKenzie et al., 2022). We note that the performance of successful runs remain monotonic during ranges of inverse scaling, confirming that U-shaped curves in performance distributions are an artifact of success probability; conventional scaling describes the performance of successful runs.
@@
  • In realistic settings, multimodal performance clusters emerge based on data composition and model scale. In Section 4, we extend our analysis to the task of multiple choice question answering by continually pretraining LLMs. Our findings confirm that random variation in LLMs is also bimodal on established emergent capabilities.
@@
Figure 2: Histograms of exact match accuracy when generalizing to length $40$ sequences on the reverse order addition task when independently scaling width (above) and depth (below).
2 Methodology
@@ 2.1 Synthetic Tasks
Breakthrough capabilities often require compositional reasoning (Srivastava et al., 2023; Löwe et al., 2024; Chen et al., 2024); we consider the compositional property of length generalization (Graves et al., 2016; Kaiser & Sutskever, 2015; Lake & Baroni, 2018; Hupkes et al., 2020). Below we outline our experimental setup, with further details in Appendix A.
Architecture: In our synthetic experiments, we train decoder-only Transformer models from scratch using rotary position embeddings (RoPE) (Su et al., 2024). To observe the random performance distribution at each scale, we train our models from hundreds of seeds. We choose model sizes by separately increasing the width and depth hyperparameters, fixing the other hyperparameter. We scale width by increasing the number of 64-parameter heads per layer.
Task: We consider two algorithmic tasks previously studied in Zhou et al. (2024a): counting and reverse-order addition. Performance on these task ranges widely across random seeds. At each scale, we train models from 250 seeds for learning to count and 200 seeds for reverse-order addition.
  • Count: Given two numbers in increasing order, the model is trained to generate a sequence which counts consecutively from the first number to the second number. Examples are given in the form "5, 9 >, 5, 6, 7, 8, 9", while limiting the length of the counting sequence during training. Zhou et al. (2024a) showed that models trained to count can generalize to more than twice this training length; however, we will reveal a more nuanced view of length generalization based on its distribution across independent model runs.
  • Reverse-Order Addition with Index Hints: Zhou et al. (2024b) show that Transformers can generalize to 10-15 digits past their training length with two modifications made to the original addition task: generating the digits of the answer in reverse order and using index hints. Examples are given in the form "a0, 3, a1, 4, +, a0, 2, a1, 8, >, a1, 2, a0, 6", where the index hints are a consecutive sequence of the max number of digits sampled randomly from 0 to the max evaluation length.
Dataset: During training, we sample sequences i.i.d from the train set and invoke in-context learning by adding examples to the context, following previous work (Jelassi et al., 2024; Zhou et al., 2024b). The length of examples are sampled uniformly from 1 to the maximum training length, set to 30 for count and 35 for reverse-order addition.
@@ 2.2 Natural Language Tasks
We extend our investigation to LLMs by testing an emergent natural language understanding task, multiple choice question answering. We test whether the performance clusters observed in emergent synthetic tasks also manifest in LLMs. Specifically, we hypothesize that emergent scaling curves in LLMs express underlying bimodal (or multimodal) performance distributions. During training, the large and diverse pretraining corpus encourages models to acquire various capabilities. While sufficiently large models may learn all such capabilities, smaller models have limited capacity, requiring these capabilities to compete (Merrill et al., 2023). This competition, influenced by initialization and data order, leads to varying outcomes across random seeds, forming performance clusters for specific benchmarks.
We train LLMs to answer multiple-choice questions, an established breakthrough capability (Srivastava et al., 2023; Snell et al., 2024). Our procedure can be regarded as continued pretraining, as we mimic the model’s original pretraining objective with a dataset of English language sequences. We hypothesize that at the emergence point for this task, performance distributions will also be bimodal across random seeds.
Task: We consider finetuning LLMs for the MMLU benchmark (Hendrycks et al., 2021), a multiple-choice question-answering task. Achieving strong performance on MMLU requires two key abilities: (1) natural language understanding and reasoning with domain-specific knowledge and (2) producing correct answers in the required format. It is this latter auxiliary ability which leads to emergent trends (Srivastava et al., 2023; Hu & Frank, 2024).
Model: We use the Qwen2.5 family of base models (Yang et al., 2024). To introduce randomness during continued pretraining, we reinitialize the final attention layer and the subsequent LLM head. We perform full-parameter continued pretraining on these reinitialized models.
Data: To ensure that the multiple choice formatting circuits compete with other tasks during continued pretraining, we create a diverse dataset by mixing the English news subset of C4 (Raffel et al., 2023) with the official MMLU auxiliary training data. The C4 news data reinforces general language modeling ability, while the MMLU data focuses on multiple-choice question answering. We vary the proportion of MMLU in this mixture to control the target task’s data size in addition to model size, as both are common control variables in the scaling laws literature.
@@ 3 Synthetic Experiment Results
Existing studies on performance breakthroughs across scales typically report results for a single model or, at most, the average of a few runs. The literature suggests that these emergent behaviors are unlocked at certain model scales (Wei et al., 2022), implying that scaling curves for different model runs would generally perform similarly to within a margin of noise. Contrary to this belief, we demonstrate that breakthroughs in compositional tasks are the result of a stochastic distribution which changes gradually even as individual scaling curves jump abruptly. We will introduce these concepts and our framework through a case study of the reverse order addition task from Section 2.1.
@@ 3.1 Emergence is a Sign of Bimodal Variation
First, we report benchmark performance across scales on our synthetic tasks by emulating the conventional approach in reporting results for single seeds. Following Srivastava et al. (2023), we take the vector of model performances for the count task given a fixed seed at test length 60 across different scales and calculate their breakthroughness and linearity metrics. As defined in Appendix B, the former metric measures emergence, whereas the latter measures a smooth response to scale. We plot the performance across scale for the seeds with the highest breakthrough and highest linearity in Figure 1, fixing the random initialization and data order seed across scales. [^1] Although the fixed seed is standard practice for reporting LLM benchmark performance across scales, the initializations produced by a single seed have no meaningful relation across different scales. As seen in Figure 1 and Appendix Figure 10, we can easily find fixed seeds that show varying levels of emergence and linearity, due to random variation in breakthroughs.
This variation is explained by the bimodality of model performance distributions when varying seed. The histograms shown in Figure 2 illustrate that, for a population of models independently trained on the reverse-order addition task, performance clusters into high and low component modes at many parameter sizes. In other words, these scales exhibit distinctly bimodal distributions in length generalization capabilities. This variability is precisely what causes some model runs to appear as breakthroughs while others follow a more linear progression. Specifically, when different runs are clustered into very high or very low performance due to their bimodality, a model might exhibit linear scaling if sampled from the same cluster as the previous scale or emergent scaling when switching from the low cluster to the high cluster. These differences ultimately lead to high variability in the timing and degree of emergence.
@@ 3.2 Sudden Jumps From Gradual Distribution Shifts
When experiments report a performance metric from only one seed or a scalar value summarizing a few seeds, the outcome is likely to be close to the mode performance of the underlying model population. As shown in Figure 3 (top left), the mode of the performance distribution shows a massive spike in improvement mirroring an emergent benchmark’s scaling trend. However, we claim that this discontinuous improvement is only an artifact of sampling the mode performance, although this artifact which will also produce discontinuity in many single-seed scaling curves. Underlying this discontinuous performance jump are continuous changes in other distributional statistics. In Figure 3 (top right), the mean exhibits a smoother trend in accuracy. Mode and mean diverge thus because the underlying distribution is bimodal, expressing a mixture of “successful” and “failing” runs.
@@
Figure 3: Summary statistics for models generalizing to test length 40 on the reverse-order addition task. We track overall mode (top left) and overall mean (top right). While the mode exhibits a sharp increase in accuracy at a certain model scale whether scaling width or depth, the mean evolves more continuously as a result of the bimodal nature of the random variation distribution. We also take note of the part of the distribution corresponding to successful runs by plotting the fraction of runs reaching above 20% accuracy (bottom left) and the mean of such runs (bottom right). The plots for mean include 95% confidence intervals with 1000 bootstrapped samples.
Figure 4: Scaling depth and width independently for the reverse-order addition models, we show Wasserstein-L2 distance from the performance distribution of the final, largest model. We mark the emergence of bimodality based on the last scale before multiple peaks appear in Figure 2. We mark the mode breakthrough at the point where the mode in Figure 3 (top left) sharply increases because successful length generalization becomes marginally more likely than failure. Note that the sharp decrease in W2 distance—corresponding to the appearance of bimodality at depth 3 in Figure 2—occurs before the mode breakthrough. During the apparent breakthrough in the mode, the distributions shift more gradually and continuously with increases in scale.
(a) Fixing network depth to 4 layers.
(b) Fixing network width to 4 heads per layer.
Figure 5: Histograms for EM accuracy for test length $60$ on the count task, across 8 different scales for independently scaling depth (above) and 10 different scales for scaling width (below).

### v3

# Distributional Scaling of Emergent Capabilities

@@ Abstract
This paper explores the nature of sudden breakthroughs in language model performance at scale, which stand in contrast to smooth improvements governed by scaling laws. While advocates of “emergence" view breakthroughs as unlocked capabilities, others attribute them to thresholding effects on noncontinuous metrics. We propose that breakthroughs are instead driven by continuous changes in the probability distribution of training outcomes when performance is bimodally distributed across random seeds. In synthetic length generalization tasks, we show that different random seeds can produce either highly linear or emergent scaling trends. We reveal that sharp breakthroughs in metrics are produced by underlying continuous changes in their distribution across seeds. Furthermore, we provide a case study of inverse scaling and show that even as the probability of a successful run declines, the average performance of a successful run continues to increase monotonically. We validate our distributional scaling framework on realistic settings by measuring MMLU performance in LM populations. These insights emphasize the role of random variation in the effect of scale on LM capabilities.
@@ 1 Introduction
On most benchmarks, language model (LM) performance is determined by a scaling law [12, 35, 21] that varies smoothly with parameter size and overall training compute. There are, however, a number of celebrated exceptions in which performance abruptly improves on specific benchmarks [41].
@@
These sudden breakthroughs fuel one of the most heated debates in modern machine learning. On one side, advocates of emergence claim that performance abruptly improves at particular scales because those scales allow the LM to acquire specific concepts that permit out-of-distribution generalization [45]. On the other side, skeptics argue that these sudden improvements are a mirage driven by thresholding effects and alleviated by more appropriate continuous metrics—though a few breakthrough capabilities remain stubbornly emergent [36]. Here, we argue that such discontinuities are driven by continuous changes in the probability of a breakthrough at each scale.
@@
We posit that a breakthrough capability is often distinguished not by deterministic responses to scale, but by multimodal random variation—in other words, independent training runs form multiple clusters in their performance metrics. Such effects are currently undetected because scaling laws are usually determined by measuring a single training run at each scale, and resources are rarely committed to correct for random variation by reusing the same hyperparameter settings with different random seeds. Although random variation may be benign when model performance is measured in-distribution [18], previous work suggests that out-of-distribution performance may vary widely across training runs [49, 51], even at larger scales [26]. Since training numerous seeds becomes prohibitively expensive at large scales, we first study length generalization in small models trained on synthetic tasks. In similar settings, existing publications consider only a few training runs and report their summary statistics. By contrast, we characterize their full distribution, demonstrating that these breakthrough tasks exhibit multimodal variation. Our findings can be summarized as follows:
@@
Figure 1: Different random seeds produce different scaling trends. Scaling trends can be emergent or linear for different seeds, even if all models train on the same data with the same hyperparameters. On the counting task, we show trends for random seeds with the highest breakthroughness (seed 93; top left) and linearity (seed 205; top right). We mark parameter counts immediately before and after seed 93’s emergence respectively as (a) and (b). Histograms illustrate the bimodal distribution of performance across all random seeds at scales (a) and (b), marking the positions of seeds 93 and 205. Breakthroughs occur when consecutive points represent different clusters; linear trends occur when each point is sampled from the same gradually shifting cluster.
@@
  • Breakthrough scaling results from bimodal performance distributions. In language modeling and synthetic length generalization case studies, we demonstrate that some random seeds produce linear scale trends while others are emergent. This variation is caused by the bimodal distribution of each skill across random seeds, a property that materializes around breakthrough thresholds in model size. At these scales, “emergence” is a stochastic property. We confirm these capabilities remain bimodally distributed under continuous metrics.
@@
  • While a scale curve exhibits sudden, discontinuous improvement in a skill, the probability of learning that skill may be responding continuously to scale. Modeling the bimodal distribution as a mix of failure and success distributions, we illustrate that average improvements can come from changes in the probability of success or in the mean performance of a successful model (Section 3). Measuring the Wassterstein distances between distributions at each scale (Section 3.3), we show that bimodality appears abruptly at some minimum capacity scale, but these sudden changes do not necessarily align with emergent trends for individual model runs.
@@
  • Competing solutions can lead to either monotonic or U-shaped trends in emergence likelihood. We discuss a synthetic task that exhibits periods of inverse scaling in average improvement (Section 3.4), a phenomenon also observed in LMs [28]. We note that the mean performance of successful runs remain monotonic during ranges of inverse scaling, suggesting that U-shaped curves in performance distributions are an artifact of success probability.
@@
  • In realistic settings, multimodal performance clusters emerge based on data composition and model scale. In Section 4, we extend our analysis to the task of multiple choice question answering by continually pretraining LMs. Our findings confirm that random variation in LMs is also bimodal for established emergent capabilities near their breakthrough threshold.
@@
Figure 2: Random variation in length generalization (addition task). Histograms of exact match accuracy on length $40$ sequences when independently scaling 2(a) width and 2(b) depth.
2 Experiments
@@ 2.1 Synthetic Tasks
We measure models’ length generalization [10, 20, 23, 15], one of many compositional properties that lead to emergence [41, 25, 3]. Experimental setup is described below, with details in Appendix A.
Architecture: In our synthetic experiments, we train decoder-only Transformer models from scratch using rotary position embeddings (RoPE) [42]. To observe the random performance distribution at each scale, we train our models from hundreds of seeds. We choose model sizes by separately increasing the width and depth hyperparameters, fixing the other hyperparameter. We scale width by increasing the number of 64-parameter heads per layer.
Task: We consider two algorithmic tasks previously studied in Zhou et al. [49]: counting and reverse order addition. Performance on these task ranges widely across random seeds. At each scale, we train models from 250 seeds for the count task and 200 seeds for reverse order addition.
  • Count: Given two numbers in increasing order, the model is trained to generate a sequence which counts consecutively from the first number to the second number. Examples are given in the form "5, 9 >, 5, 6, 7, 8, 9", while limiting the length of the counting sequence during training. Zhou et al. [49] showed that models trained to count can generalize to more than twice this training length; however, we will reveal a more nuanced view of length generalization based on its distribution across independent model runs.
  • Reverse Order Addition with Index Hints: Zhou et al. [51] showed that Transformers can generalize 10-15 digits past training length for an addition task, if provided with index hints and allowed to generate the answer backwards. Examples take the form "a0, 3, a1, 4, +, a0, 2, a1, 8, >, a1, 2, a0, 6", where index hints are a consecutive sequence of the max number of digits sampled randomly from 0 to the max evaluation length. In this work, we refer to this task as the ‘addition task’ despite these modifications.
Dataset: During training, we sample sequences i.i.d from the train set and invoke in-context learning by adding examples to the context, following previous work [17, 51]. The length of examples are sampled uniformly from 1 to the maximum training length, set to 30 for count and 35 for addition. Length generalization is then tested at length 60 for count and 40 for addition.
@@ 2.2 Natural Language Tasks
We extend our investigation to LMs by testing multiple choice question answering, an emergent natural language understanding task [41, 40]. (A second task, CommonSenseQA, is in Appendix E.) We test whether the performance clusters observed in emergent synthetic tasks also manifest in LMs. Specifically, we hypothesize that emergent scaling curves in LMs express underlying bimodal (or multimodal) performance distributions.
Our procedure can be regarded as continued pretraining, as we mimic the model’s original pretraining objective with a dataset of English language sequences. We hypothesize that at the emergence point for this task, performance distributions will also be bimodal across random seeds. During pretraining, a large diverse corpus encourages models to acquire various capabilities. While sufficiently large models may learn all such capabilities, smaller models have limited capacity, requiring these capabilities to compete [29]. This competition, influenced by initialization and data order, leads to varying outcomes across random seeds, forming performance clusters for specific benchmarks.
Task: We consider finetuning LMs for the MMLU benchmark [11], a multiple-choice question-answering task. Achieving strong performance on MMLU requires two key abilities: (1) natural language reasoning with domain-specific knowledge and (2) producing correct answers in the required format. It is this latter ability which leads to emergent trends [41, 13].
Model: We use the Qwen2.5 family of base models [47]. To introduce randomness during continued pretraining, we reinitialize the final attention layer and the subsequent LM head. We perform full-parameter continued pretraining on these reinitialized models.
Data: To ensure that the multiple choice formatting circuits compete with other tasks during continued pretraining, we create a diverse dataset by mixing the English news subset of C4 [34] with the official MMLU auxiliary training data. The C4 news data reinforces general language modeling ability, while the MMLU data focuses on multiple-choice question answering. We vary the proportion of MMLU in this mixture to control the target task’s data size in addition to model size, as both are common control variables in the scaling laws literature.
@@ 3 Synthetic Experiment Results
Existing studies on performance breakthroughs across scales typically report results for a single model or, at most, the average of a few runs. The literature suggests that these emergent behaviors are unlocked at certain model scales [45], implying that scaling curves for different model runs would generally perform similarly to within a margin of noise. Contrary to this belief, we demonstrate that breakthroughs in compositional tasks are the result of a stochastic distribution which changes gradually even as individual scaling curves jump abruptly. We will introduce these concepts and our framework through a case study of the addition task from Section 2.1.
@@ 3.1 Emergence is a Sign of Bimodal Variation
First, we report benchmark performance across scales on our synthetic tasks by emulating the conventional approach in reporting results for single seeds. Following Srivastava et al. [41], we take the vector of model performances for the count task given a fixed seed at test length 60 across different scales and calculate their breakthroughness and linearity metrics. As defined in Appendix B, the former metric measures emergence, whereas the latter measures a smooth response to scale. We plot the performance across scale for the seeds with the highest breakthrough and highest linearity in Figure 1, fixing the random initialization and data order seed across scales. [^1] Although it is standard practice to fix the random seed when reporting LM benchmark performance across scales, the initializations produced by a single seed have no meaningful relation across different scales. As seen in Figure 1 and Appendix Figure 8, we can easily find fixed seeds that show varying levels of emergence and linearity, due to random variation in breakthroughs.
This variation is explained by the bimodality of model performance distributions when varying seed. The histograms shown in Figure 2 illustrate that, for a population of models independently trained on the addition task, performance clusters into high and low component modes at many parameter sizes. In other words, these scales exhibit distinctly bimodal distributions in length generalization capabilities. This variability is precisely what causes some model runs to appear as breakthroughs while others follow a more linear progression. Specifically, when different runs are clustered into very high or very low performance due to their bimodality, a model might exhibit linear scaling if sampled from the same cluster as the previous scale or emergent scaling when switching from the low cluster to the high cluster. These differences ultimately lead to high variability in the timing and degree of emergence.
@@ 3.2 Sudden Jumps From Gradual Distribution Shifts
When reporting metrics from only one seed or the mean of a few seeds, the outcome is likely to be close to the mode performance of the underlying model population. As shown in Figure 3 (top left), the mode of the performance distribution shows a massive spike in improvement mirroring an emergent benchmark’s scaling trend. However, we claim that this discontinuous improvement is only an artifact of sampling the mode performance, although this artifact which will also produce discontinuity in many single-seed scaling curves. Underlying this discontinuous performance jump are continuous changes in other distributional statistics. In Figure 3 (top right), the mean exhibits a smoother trend in accuracy. Mode and mean diverge thus because the underlying distribution is bimodal, expressing a mixture of “successful” and “failing” runs.
@@
Figure 3: Summary statistics for length generalization (addition task). Exact match statistics for 200 models trained at length 35 and tested at length 40. We track overall 3 mode and 3 mean. Because the EM accuracy distribution is bimodal, the mode exhibits a sharp increase even as the mean evolves continuously. Defining success as >20% EM accuracy, we also note the continuous change in 3 the probability of success and in 3 the mean of successful runs . Means feature 95% confidence intervals over 1000 bootstrapped samples.
(a) Addition task. The sharpest change in distance occurs early. During the mode breakthrough, the distributions are still shifting gradually.
(b) Count task. The distribution changes slowly between intermediate model scales, but changes suddenly at the nadir of the U-shaped curve in Figure 5.
Figure 4: Changes in random variation. Wasserstein-L2 distance of each scale’s performance distribution relative to the largest scale, scaling depth and width independently. We mark the emergence of bimodality at the last scale before multiple peaks appear. We mark the mode breakthrough at the last scale before successful length generalization becomes marginally more likely than failure.